Can someone help me on this problem?
I need to solve for x and check my results:
6x-5/5 - 2x/3 = 3
Thanks,
Sher

Found 2 solutions by stanbon, checkley71:
Answer by stanbon(75887)

(6x-5)/5 - 2x/3 = 3
Multiply thru by 15.
3(6x-5) - 10x = 45
18x-15-10x = 45
8x=60
x=15/2
Cheers,
Stan H.

Answer by checkley71(8403)

6X-5/5-2X/3=3 OR 6X-1-2X/3=3 OR 6X-2X/3=3+1 OR (18X-2X)/3=4 OR 16X/3=4 OR
16X=12 OR X=12/16 OR X=3/4
PROOF 6(3/4)-1-2(3/4)/3=3 OR 18/4-1-(6/4)/3=3 OR 18/4-1-(1.5)/3=3 OR 4.5-1-.5=3
OR 3=3